OpenStudy (anonymous):

(0,3) slope= -1/4 Find the slope intercept and standard form

OpenStudy (anonymous):

slope intercept= y=mx+b Standard Form= AX+BY=C So: 3=...?

OpenStudy (anonymous):

m is the slope, right?

OpenStudy (ivettef365):

right, and b is the y-int

OpenStudy (anonymous):

how would i find the y intercept?

OpenStudy (ivettef365):

is given point (0,3), so what is the y when x =0

OpenStudy (anonymous):

y=3

OpenStudy (ivettef365):

so b = 3, so what is the slope intercept formula

OpenStudy (anonymous):

3=-1/4(0)+3 thanks so much!

OpenStudy (anonymous):

but i don't get what numbers to plug in to get standard form.

OpenStudy (ivettef365):

when you write the slope intercept, you write it like this y = -1/4x+3 that is the equation with slope -1/4 and y-int 3 the x and y could be any points on that line so you don't put values

OpenStudy (anonymous):

thank you! but how would i find standard form?

OpenStudy (ivettef365):

when you have the equation in slope form you can change it on that website

OpenStudy (anonymous):

wow! thank so much!

OpenStudy (ivettef365):

let me know if you know how to do it they give you a window like this y = ____ x + ____

OpenStudy (anonymous):

ok. one moment

OpenStudy (ivettef365):

you enter -1/4 on the first blank and 3 on the second

OpenStudy (anonymous):

i got 1x+4x=12

OpenStudy (anonymous):

thank you!

OpenStudy (ivettef365):

good job

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

oops :)

OpenStudy (ivettef365):

but it should b 1x + 4y = 12

OpenStudy (anonymous):

yeah. thanks :)

OpenStudy (ivettef365):

ok, good